A solution contains 6.56x103 M silver acetate and 1.28x102 M lead nitrate. Solid ammonium iodide is added slowly to this mixture. A. What is the formula of the substance that precipitates first? formula B. What is the concentration of iodide ion when this precipitation first begins? [iodide] = |
Answer #1

ammonium iodide is NH4I

iodide ions will form

Ag+ + I- = AgI

Pb+2 + I- = PbI2

Get Ksp data:

AgI 8.52×10–17; PbI2 9.8×10–9

Therefore, Ksp << Ag than PbI2

then, most likely AgI precipitates first

b)

find [I-] when it begins to precipitate

Ksp = [Ag+][I-]

8.52*10^-17 = (6.56*10^-3)[I-]

[I-] = (8.52*10^-17) / (6.56*10^-3) = 1.2987*10^-14 M